Circular No. 6520

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bw IN NGC 664 S. Benetti and M. Turatto, European Southern Observatory, report:

“A fully-reduced CCD spectrogram of SN 1996bw (range 370–970 nm, resolution 1.5 nm) obtained on Dec. 7.2 UT with the 3.6-m telescope (+ EFOSC) shows the typical spectral features of a type II supernova about two weeks after maximum. The spectrum consists of a broad Hα emission that is now developing a broad P-Cyg absorption, and there are welldeveloped P-Cyg profiles of Hδ, Hγ, Hβ and He I 587.6 nm. Blueward of Hα an absorption identified with Si II 635.5 nm is also present. A similar feature was visible in the spectra of SNe 1990E and 1992H ∼ 19 days after maximum (Clocchiatti et al.) The lines are superimposed on a blue continuum (T_{bb} ∼ 8600 K). The expansion velocity deduced from the minima of the lines is ∼ 11 000 km/s.”

COMET 126P/IRAS

Continuation to the ephemeris on IAUC 6483:

1996 /97 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 3 22^{h}05.^{m}49 + 2^{o}07.^{′}9 1.559 1.743 83.^{o}2 34.^{o}2 12.8
13 22 17.70 + 6 43.7 1.689 1.770 78.0 33.0 13.1
23 22 31.22 +10 58.9 1.819 1.802 73.3 31.5 13.4
Jan. 2 22 45.86 +14 58.4 1.950 1.840 68.9 29.9 13.7
12 23 01.53 +18 45.9 2.078 1.883 64.8 28.2 14.1
22 23 18.16 +22 24.1 2.204 1.930 61.0 26.5 14.4
Feb. 1 23 35.72 +25 54.5 2.328 1.982 57.5 24.8 14.8
11 23 54.23 +29 18.0 2.448 2.036 54.3 23.2 15.1
21 0 13.69 +32 35.0 2.565 2.094 51.3 21.6 15.5
Mar. 3 0 34.17 +35 45.1 2.679 2.154 48.5 20.2 15.8
13 0 55.72 +38 47.6 2.789 2.217 46.0 18.8 16.1

Total visual magnitude estimates (L = reflector): Nov. 1.89 UT, 11.8 (J. Carvajal, Teruel, Spain, 0.44-m L); 8.77, 12.6 (K. Sárneczky, Ráktanya, Hungary, 0.44-m L); 11.19, 12.6 (A. Hale, Cloudcroft, NM, 0.41-m L); 29.79, 12.8 (D. Rodriguez, Toledo, Spain, 0.44-m L); Dec. 1.75, 12.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m L); 3.74, 12.8 (Hornoch).

CORRIGENDA

IAUC 6465, Comet C/1995 O1 (Hale-Bopp), for 17^{h}23^{m} read 17^{h}29^{m} IAUC 6518, γ-Ray Bursts, for 1^{o}.4, 3^{o}.0 and 1^{o}.0, read 1.4, 3.0 and 1.0

standard deviations

1996 December 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden